OpenStudy (anonymous):

Which graph represents the function of f(x) = 4x^-16/2x-4 ?

OpenStudy (anonymous):

a.

OpenStudy (anonymous):

b.

OpenStudy (anonymous):

c.

OpenStudy (anonymous):

or d.

OpenStudy (campbell_st):

just factor the numerator and denominator... remove common factors and you get a straight line. find the intercepts and you'll have your line.

OpenStudy (only4u):

4x^ ?????

OpenStudy (anonymous):

also. before you factor you have to find the restriction on x. so you know where the graph is discontinuous

OpenStudy (anonymous):

which one is it?

OpenStudy (anonymous):

\[\frac{4x^2-16}{2x-4}=\frac{4(x^2-4)}{2(x-2)}=\frac{2(x+2)(x-2)}{x-2}=2(x+2)=2x+4\] is a start

OpenStudy (anonymous):

your answer is whatever line has slope \(2\) and \(y\)-intercept \(4\)

OpenStudy (anonymous):

oh i see, you have to such lines, pick the one with the hole in it at \((2,8)\) because \(x\) cannot be \(2\)

OpenStudy (anonymous):

is it c then?